a chemist mixes 150ml of solution that is 45% acid with 100ml of pure acid
how many ml of acid are in the resulting mixture?
what percentage of the resulting mixture is acid

Found 2 solutions by addingup, stanbon:
Answer by addingup(3677)

0.45(150)+100 = 167.5
So the total of the new mixture will be 150+100 = 250ml of which 167.5 are acid
167.5/250 = 0.67 so 67% of the resulting mixture is acid

Answer by stanbon(75887)

A chemist mixes 150ml of solution that is 45% acid with 100ml of pure acid
how many ml of acid are in the resulting mixture?
what percentage of the resulting mixture is acid.
---------
Equation:
acid + acid = acid
0.45*150 + 1*100 = 167.5 ml of acid
------
Percentage = 167.5/(150+100) = 0.67 = 67%
----------
Cheers,
Stan H.
--------